given a rectangle has a length of x+2 and a width of x, write an expression in simplified form for the perimator of the rectangle
Answer by nerdybill(7384)

Perimeter = 2(length + width)
.
Since:
length = x+2
width = x
we now have:
Perimeter = 2(x+2 + x)
Perimeter = 2(2x+2)
Perimeter = 4x+4
.
Answer: 4x+4